# Short-Range Order Based Ultra Fast Large-Scale Modeling of High-Entropy Alloys

### v2

@@ Abstract
High-Entropy Alloys (HEAs) exhibit complex atomic interactions, with short-range order (SRO) playing a critical role in determining their properties. Traditional methods, such as Monte Carlo (MC) based ATAT and SCRAPs, and hybrid Monte Carlo-Molecular Dynamics (MC-MD), are often constrained by system size and computational inefficiency, limiting their applicability for large-scale modeling. We introduce PyHEA, a Python based toolkit with a high-performance C++ core, designed to efficiently model HEAs through a combination of global and local search algorithms, incremental SRO calculations, and GPU acceleration. PyHEA delivers an impressive speedup of more than 333,000× and 10,000× over state-of-the-art tools ATAT and SCRAPs respectively, ensuring accurate SRO calculations with unprecedented efficiency. In practical applications, PyHEA successfully modeled a 256,000-atom Fe-Mn-Cr-Co alloy, replicating literature results and reducing computational time from days to seconds with GPU acceleration, achieving over 1,000× speedup compared to hybrid MC-MD approaches. This huge breakthrough in computational efficiency revolutionizes HEA modeling by bridging theoretical predictions and practical applications, opening new horizons for the design of next-generation HEAs.
@@
[1]organization=School of Software and Microelectronics, Peking University, city=Beijing, country=China
@@
[2]organization=School of Integrated Circuits, Peking University, city=Beijing, country=China
@@ 1 Introduction
High-Entropy Alloys (HEAs), introduced by Jien-Wei Yeh et al. in 2004 [1], have emerged as a transformative class in materials science. Characterized by the presence of five or more principal elements in quasi-equimolar ratios (typically 3 to 55% atomic percent each), HEAs exhibit a unique combination of remarkable properties, including high strength, hardness and exceptional resistance to oxidation and corrosion [2, 3, 4, 5]. These attributes position HEAs as promising materials for applications in extreme environments, such as aerospace, nuclear, and energy sectors. Furthermore, the unconventional nature of HEAs challenges traditional alloy design theories, which spurred significant research interest in their underlying principles [6, 7].
Recent advancements have highlighted the critical role of microstructure in determining HEAs’ extraordinary properties, particularly the influence of short-range order (SRO) [8, 9]. SRO refers to the local atomic arrangement within an alloy, a phenomenon that becomes increasingly prominent in complex concentrated alloys, even under rapid solidification conditions [11, 13]. The impact of SRO on mechanical performance and thermal stability is profound, influencing elastic modulus, stiffness, and ductility by optimizing the balance between strength and toughness [14, 15, 16]. These findings underscore the potential for leveraging SRO control to tailor HEA properties, providing a theoretical framework for the development of new high-performance materials.
@@
Current approaches to HEA modeling face significant limitations in addressing this challenge. Static structure construction methods, such as ATAT [21, 42, 22] and SCRAPs [23], rely on Special Quasirandom Structures (SQS) and hybrid Monte Carlo approaches, respectively, to generate random alloy configurations. While effective for handling relatively small supercells (tens to hundreds of atoms), these methods become computationally prohibitive for larger systems [20] and fail to capture critical SRO effects, particularly at lower temperatures where atomic ordering significantly influences material properties [24, 25, 26].
Hybrid Monte Carlo-Molecular Dynamics (MC-MD) methods offer an alternative approach by incorporating MC steps into extended MD simulations. Although hybrid MC-MD can theoretically handle systems with hundreds of thousands of atoms [27, 17, 28], its practical effectiveness depends heavily on the availability of highly accurate interatomic potentials. Even when such potentials exist, generating configurations that reliably reflect SRO demands substantial computational resources, often requiring hours or days of calculation time.
@@
    Enhanced accuracy and versatility: PyHEA significantly outperforms existing methods like ATAT and SCRAPs in both modeling efficiency and stability. Our benchmarks in Table 1 demonstrate up to 333,000 and 13,900 times speedup compared with ATAT and SCRAPs respectively while maintaining or improving accuracy. Unlike current tools that primarily generate random configurations, PyHEA supports user-defined target SRO values, making it uniquely suitable for modeling temperature-dependent ordering effects and other practical scenarios where specific atomic arrangements are desired.
@@
    Unprecedented scalability and speed: Through advanced GPU acceleration (CUDA) and optimized search algorithms, PyHEA efficiently handles systems with hundreds of thousands of atoms. For example, it successfully modeled a 256,000-atom Fe-Mn-Cr-Co system in under 40 seconds—a task that typically requires hours or days using traditional hybrid MC-MD method. The details of the hybrid MC-MD simulations are elaborated in section 4. This breakthrough in computational efficiency enables the exploration of mechanical properties and structural characteristics at previously inaccessible scales.
@@
    Comprehensive software ecosystem: Built on Python with a high-performance C++ core, PyHEA combines accessibility with computational efficiency. Released under the LGPL-3.0 license, it provides an end-to-end solution including structure generation, SRO optimization, and analysis tools. The modular design facilitates easy integration with existing workflows and enables rapid customization for specific research needs.
@@ 2.1 Parameter Optimization Analysis
Figure 1: Influence of input parameters on fitness and execution time. (a) Number of solutions, (b) global iterations, (c) number of tasks, and (d) search depth. The plots illustrate the impact of varying each parameter on the fitness (left axis) and execution time (right axis).
In optimizing atomic configurations for HEAs, several key parameters significantly influence both the quality of results and computational efficiency. These parameters include: the number of solutions (which determines the population size for exploring different atomic arrangements), global iterations (controlling how many times the entire optimization process is repeated), number of tasks (affecting the degree of parallelization in the search process), and search depth (determining how thoroughly each local configuration is explored). Each parameter plays a distinct role in the search algorithm detailed in subsection 4.1 and subsection 4.2. Understanding their optimal values is crucial for achieving the best balance between computational efficiency and solution quality.
To evaluate optimization quality, we employ a fitness function that quantifies how well atomic configurations align with the target Short-Range Order (SRO) characteristics. Lower fitness values indicate better alignment with desired SRO properties. Our analysis reveals distinct patterns in how each parameter affects both fitness and computational efficiency.
As depicted in Fig. 1(a-b), an increase in the number of solutions improves fitness; however, beyond 64 solutions, this enhancement yields diminishing returns while causing a marked rise in execution time. A similar pattern emerges for global iterations, where substantial fitness improvements are observed up to 6 iterations. Beyond this threshold, further gains are minimal, whereas execution time continues to escalate. These results indicate that using 64 to 128 solutions and 6 global iterations achieves an effective balance between precision and computational cost.
Parallelism also plays a critical role in the optimization process. As demonstrated in Fig. 1(c), increasing the number of tasks enhances the consistency of fitness outcomes. Nevertheless, execution time grows significantly after 128 tasks, with only marginal improvements in fitness. Therefore, a task count between 64 and 128 represents an optimal range. Similarly, Fig. 1(d) shows search depth affects the thoroughness of exploration within each iteration. Deeper searches (greater than 8) yield better fitness values, they also come at the expense of increased computational demands. Based on these findings, a search depth of 8-10 strikes an optimal balance between precision and computational efficiency.
Consequently, we recommend two parameter configurations optimized for different priorities. The precision-optimized configuration (128 solutions, 10 global iterations, 256 tasks, search depth 10) achieves maximum accuracy in SRO calculations. The performance-optimized configuration (64 solutions, 6 global iterations, 64 tasks, search depth 8) maintains good optimization quality while reducing computation time as much as possible. Users can select either configuration based on their specific requirements for precision versus computational efficiency.
@@ 2.2 Comparison with State-of-the-Art Methods
To evaluate the performance of PyHEA, we conducted a comparative analysis against SCRAPs [23], a state-of-the-art method commonly utilized in HEA modeling. This comparison focused on evaluating both fitness quality and computational efficiency across varying system sizes.
@@
To further validate PyHEA’s acceleration benefits, we also conducted a comparison with another widely used MC-based tool, ATAT. Table 1 summarizes the performance metrics based on serial execution across PyHEA (CPU and GPU), SCRAPs, and ATAT. While SCRAPs and ATAT perform comparably for smaller system sizes, they both struggle as the number of atoms increases. Like SCRAPs, ATAT becomes impractical beyond a few hundred atoms. Conversely, PyHEA’s GPU implementation maintains substantially lower execution times across all tested system sizes, underscoring its performance advantages and scalability. For a comprehensive comparison of parallel CPU performance between these tools, readers are referred to Table 2 in the Appendix.
Table 1: Performance comparison between PyHEA and existing methods (ATAT and SCRAPs) for optimizing atomic structures in high-entropy alloys, focusing on serial execution performance (parallel performance data available in Appendix). The table shows computation times (in minutes) and speedup ratios for both body-centered cubic (bcc) and face-centered cubic (fcc) structures across varying system sizes (#At) and numbers of species (#Sp). Both PyHEA and SCRAPs were executed in serial mode using a single CPU core (Intel Xeon Gold 6132 @ 2.60GHz), with additional GPU (NVIDIA 4090) results shown for PyHEA. The optimizations used 128 candidate solutions over 10 global iterations with 256 parallel MC tasks (search depth 10). Supercell sizes followed $S=A\cdot L^{3}$, where $A=2$ for bcc or $A=4$ for fcc structures. All calculations optimized short-range order (SRO) over 3 coordination shells. Dashes (–) indicate computationally infeasible cases, while speedup values are shown relative to SCRAPs[ATAT in brackets].
  Type #Sp #At ATAT SCRAPs PyHEA PyHEA Speedup
Serial Serial GPU CPU GPU
@@
    ATAT data from Ref. [23]. All times in minutes.
@@ 2.3 Application
Figure 3: Combined visualization of SRO and radial distribution functions (RDF) for atomic pair interactions at different temperatures. (a) The radar plot illustrates the SRO values for key atomic pairs (e.g., Fe-Fe, Fe-Mn, Fe-Cr) at three temperatures: 750 K, 875 K, and 1373 K. (b) The RDF plots display the spatial distribution of atomic pairs as a function of separation distance (Å) for the same temperatures, with smooth curves representing each pair.
To illustrate the practical potential of PyHEA, we conducted large-scale molecular dynamics (MD) simulations using LAMMPS, modeling a system containing 256,000 atoms of Fe-Mn-Cr-Co (composition: 40% Fe, 40% Mn, 10% Cr, and 10% Co) at various temperatures. The system was generated using PyHEA, enabling us to bypass the conventional Monte Carlo (MC) and MD hybrid process and achieve significant reductions in simulation time.
The system was simulated in an NPT ensemble for 10 ps, with the Radial Distribution Function (RDF) calculated to analyze the SRO behavior at different temperatures. As depicted in Fig. 3, our results align closely with those from a reference study in which MC simulations were performed on a comparable Fe-Mn-Cr-Co system to investigate compositionally-driven SRO [33]. Additionally, our simulation outcomes were consistent with the ordering patterns reported in experimental studies [34, 35] in a $\text{Fe}_{50}\text{Mn}_{30}\text{Cr}_{10}\text{Co}_{10}$ system.
Employing PyHEA to generate input structures for LAMMPS allowed us to circumvent the extensive MC-MD hybrid computations typically required, which can take several days to complete. PyHEA reduced this computation time to just tens of seconds, showcasing a remarkable improvement in computational efficiency. Furthermore, the scale of the simulation, involving 256,000 atoms, surpasses the capabilities of most existing methods, emphasizing PyHEA’s robustness and effectiveness in handling large-scale atomistic simulations.
The temperature-dependent evolution of SRO in this system is further illustrated by Fig. 6 in the Appendix, which presents Warren-Cowley SRO parameter matrices at different temperatures (1373 K, 873 K, and 750 K). These heat maps reveal how the atomic ordering patterns shift from a nearly random distribution at high temperatures (1373 K) to more pronounced local ordering at lower temperatures (750 K), providing valuable insight into the system’s structural evolution.

### v3

@@ Abstract
High-Entropy Alloys (HEAs) exhibit complex atomic interactions, with short-range order (SRO) playing a critical role in determining their properties. Traditional methods, such as Monte Carlo generator of Special Quasirandom Structures within the Alloy Theoretic Automated Toolkit (ATAT-mcsqs), Super-Cell Random APproximates (SCRAPs), and hybrid Monte Carlo-Molecular Dynamics (MC-MD)—are often hindered by limited system sizes and high computational costs. In response, we introduce PyHEA, a Python-based toolkit with a high-performance C++ core that leverages global and local search algorithms, incremental SRO computations, and GPU acceleration for unprecedented efficiency. When constructing random HEAs, PyHEA achieves speedups exceeding 333,000× and 13,900× over ATAT-mcsqs and SCRAPs, respectively, while maintaining high accuracy. PyHEA also offers a flexible workflow that allows users to incorporate target SRO values from external simulations (e.g., LAMMPS or density functional theory (DFT)), thereby enabling more realistic and customizable HEA models. As a proof of concept, PyHEA successfully replicated literature results for a 256,000-atom Fe–Mn–Cr–Co system within minutes—an order-of-magnitude improvement over hybrid MC–MD approaches. This dramatic acceleration opens new possibilities for bridging theoretical insights and practical applications, paving the way for the efficient design of next-generation HEAs.
@@
[1]organization=School of Software and Microelectronics, Peking University, city=Beijing, postcode=100871, country=China
@@
[2]organization=School of Integrated Circuits, Peking University, city=Beijing, postcode=100871, country=China
{highlights}
High Performance: PyHEA delivers unprecedented speedups of 333,000× over ATAT-mcsqs and 13,900× over SCRAPs in random structure generation, achieving superior efficiency without compromising accuracy.
Large-Scale Compatibility: PyHEA models systems with over 256,000 atoms in seconds, enabling simulations at scales relevant to experimental studies and unlocking new opportunities for high-entropy alloy research.
GPU Acceleration: PyHEA’s CUDA-based parallel Monte Carlo implementation provides over 1,000× speedup compared to CPU-based calculations, dramatically reducing simulation times for large-scale configurations.
User-Friendly Python Ecosystem: Built on a Python interface with a high-performance C++ core, PyHEA provides an accessible yet powerful solution for structure generation, SRO analysis, and data visualization, with seamless integration into existing workflows.
@@ 1 Introduction
High-Entropy Alloys (HEAs), introduced by Jien-Wei Yeh et al. in 2004 [1], have emerged as a transformative class in materials science. Characterized by the presence of five or more principal elements in quasi-equimolar ratios (typically 3-55% atomic percent each), HEAs exhibit a unique combination of remarkable properties, including high strength, hardness, and exceptional resistance to oxidation and corrosion [2, 3, 4, 5]. These attributes position HEAs as promising materials for applications in extreme environments, such as aerospace, nuclear, and energy sectors. Furthermore, the unconventional nature of HEAs challenges traditional alloy design theories, which spurred significant research interest in their underlying principles [6, 7].
Recent advancements have highlighted the critical role of microstructure in determining HEAs’ extraordinary properties, particularly the influence of short-range order (SRO) [8, 9, 24, 25, 26]. SRO refers to the local atomic arrangement within an alloy, a phenomenon that becomes increasingly prominent in complex concentrated alloys, even under rapid solidification conditions [11, 13]. The impact of SRO on mechanical performance and thermal stability is profound, influencing elastic modulus, stiffness, and ductility by optimizing the balance between strength and toughness [14, 15, 16]. These findings underscore the potential for leveraging SRO control to tailor HEA properties, providing a theoretical framework for the development of new high-performance materials.
@@
Current approaches to HEA modeling face significant limitations in addressing this challenge. Static structure construction methods like the Monte Carlo generator of Special Quasirandom Structures within the Alloy Theoretic Automated Toolkit (ATAT-mcsqs) [21, 42, 22] and Super-Cell Random APproximates (SCRAPs) [23] rely on Special Quasirandom Structures (SQS) and hybrid Monte Carlo approaches to generate random alloy configurations respectively. While these methods are effective for handling relatively small supercells (tens to hundreds of atoms), they become computationally prohibitive when dealing with larger systems. [20]
In parallel, Hybrid Monte Carlo-Molecular Dynamics (MC-MD) methods offer another alternative approach by incorporating MC steps into extended MD simulations. Although hybrid MC-MD can theoretically handle systems with hundreds of thousands of atoms [27, 17, 28], its practical effectiveness depends heavily on the availability of highly accurate interatomic potentials. Even when such potentials exist, generating configurations that reliably reflect SRO demands substantial computational resources, often requiring hours or days of calculation time.
@@
    Enhanced efficiency: We benchmarked PyHEA against ATAT-mcsqs and SCRAPs by generating a series of random structures. As summarized in Table 1, PyHEA achieves speedups of up to $3.33\times 10^{5}$ over ATAT-mcsqs and $1.39\times 10^{4}$ over SCRAPs, all without sacrificing the accuracy of the generated configurations. These results highlight PyHEA’s robust numerical stability and superior modeling efficiency, establishing it as a powerful alternative to existing methods.
@@
    Unprecedented scalability: Beyond random structures, PyHEA allows users to specify target SRO values derived from experiments or prior simulations, enabling the modeling of temperature-dependent ordering and other specialized scenarios. When such SRO data are unavailable or difficult to obtain, as introduced in  subsection 2.3, we propose an integrated workflow that employs smaller-scale ab initio (DFT) or LAMMPS simulations to compute the desired SRO. PyHEA then leverages these inputs to construct large-scale alloys, effectively streamlining conventional MC–MD processes. For instance, in modeling a 256,000-atom Fe–Mn–Cr–Co system, PyHEA provides a 1000× speedup in configuration generation. Even accounting for the time required to prepare target SRO values, an overall improvement of at least an order of magnitude is retained, making large-scale HEA simulations far more tractable.
@@
    User-friendly software ecosystem: Built on Python with a high performance C++ core, PyHEA combines accessibility with computational efficiency. Released under the LGPL-3.0 license, it provides an end-to-end solution including structure generation, SRO analysis, and data visualization. The modular design facilitates easy integration with existing workflows and enables rapid customization for specific research needs.
@@ 2.1 Parameter Optimization Analysis
Figure 1: Influence of input parameters on fitness (left axis) and execution time (right axis): (a) Number of solutions, (b) global iterations, (c) number of tasks, and (d) search depth.
In this subsection, we examine how four key parameters influence both the optimization quality and computational efficiency when constructing atomic configurations for HEAs: 1) Number of solutions, 2) Global iterations, 3) Number of tasks, and 4) Search depth. These parameters govern the search algorithm outlined in subsection 4.1 and subsection 4.2, while the fitness function used to evaluate candidate configurations is defined in subsection 4.4 (see Equation 3).
Conceptually, the fitness function serves as a “loss function” in an optimization context: it quantifies how “far” a candidate atomic configuration is from the target SRO. Lower fitness values indicate closer alignment with the desired SRO. By iteratively minimizing this fitness, the algorithm refines candidate configurations to achieve the prescribed atomic ordering.
The first parameter, number of solutions, refers to the total number of candidate configurations (atomic arrangements) evaluated in parallel during each global search iteration. As illustrated in Fig. 1(a), increasing the number of solutions from a smaller baseline up to around 64 significantly reduces the fitness, indicating better alignment with the target SRO. Beyond 64 solutions, additional gains in accuracy become negligible, while execution time increases sharply. Therefore, a range of 64–128 solutions strikes a balance between accuracy and computational efficiency.
The second parameter, global iterations, specifies the total number of optimization cycles. Fig. 1(b) shows that fitness improves substantially up to about six iterations. Beyond this point, further improvements are marginal compared to the additional runtime. Hence, six global iterations suffice in most cases to achieve robust SRO alignment without incurring excessive computational costs.
Parallelism, governed by the number of tasks, is the third parameter under consideration. As shown in Fig. 1(c), increasing the number of tasks up to around 128 enhances the consistency of fitness outcomes. However, execution time grows considerably beyond 128 tasks, while fitness gains diminish. A task count between 64 and 128 is recommended to maintain an effective balance between throughput and result fidelity.
The final parameter, search depth, determines the thoroughness of local refinements during each iteration. Fig. 1(d) demonstrates that deeper searches (greater than a depth of 8) lead to lower fitness values, signifying more accurate SRO alignment. However, deeper searches also incur higher computational costs. Consequently, a search depth of approximately 8–10 provides an optimal trade-off between solution precision and runtime.
Building on these insights, we propose two configurations tailored to different priorities:
  • 1.
    Precision-Optimized Configuration: (128 solutions, 10 global iterations, 256 tasks, search depth 10). This setup achieves the highest fidelity to the target SRO but requires longer execution times.
  • 2.
    Performance-Optimized Configuration: (64 solutions, 6 global iterations, 64 tasks, search depth 8). This arrangement maintains good SRO alignment while minimizing runtime, yielding a practical balance between accuracy and computational efficiency.
By choosing one of these configurations—or adjusting parameters within the suggested ranges—users can tailor the optimization process to either prioritize maximum accuracy or expedite computations, depending on the demands of their particular HEA simulations.
@@ 2.2 Comparison with State-of-the-Art Methods
To evaluate the performance of PyHEA, we conducted a comparative analysis against SCRAPs, a state-of-the-art method in HEA modeling[23]. This comparison focused on evaluating both fitness quality and computational efficiency across varying system sizes.
@@
To further validate PyHEA’s acceleration benefits, we also conducted a comparison with another widely used MC-based tool, ATAT-mcsqs. Table 1 summarizes the performance metrics based on serial execution across PyHEA (CPU and GPU), SCRAPs, and ATAT-mcsqs. While SCRAPs and ATAT-mcsqs perform comparably for smaller system sizes, they both struggle as the number of atoms increases. Like SCRAPs, ATAT-mcsqs becomes impractical beyond a few hundred atoms. Conversely, PyHEA’s GPU implementation maintains substantially lower execution times across all tested system sizes, underscoring its performance advantages and scalability. For a comprehensive comparison of parallel CPU performance between these tools, readers are referred to Table 2 in the Appendix.
Table 1: Performance comparison between PyHEA and existing methods (ATAT-mcsqs and SCRAPs) for optimizing atomic structures in HEAs, focusing on serial execution performance (parallel performance data available in Appendix). The table shows computation times (in minutes) and speedup ratios for both body-centered cubic (bcc) and face-centered cubic (fcc) structures across varying system sizes (#At) and numbers of species (#Sp). Both PyHEA and SCRAPs were executed in serial mode using a single CPU core (Intel Xeon Gold 6132 @ 2.60GHz), with additional GPU (NVIDIA 4090) results shown for PyHEA. The optimizations used 128 candidate solutions, 10 global iterations, 256 parallel MC tasks, and a search depth of 10. Supercell sizes followed $S=A\cdot L^{3}$, where $A=2$ for bcc or $A=4$ for fcc structures. All calculations optimized SRO over 3 coordination shells. Dashes (–) indicate computationally infeasible cases, while speedup values are shown relative to SCRAPs[ATAT in brackets].
  Type #Sp #At ATAT- SCRAPs PyHEA PyHEA Speedup
mcsqs Serial Serial GPU CPU GPU
@@
    ATAT-mcsqs data from Ref. [23]. All times in minutes.
@@ 2.3 Application
Figure 3: Schematic illustration of the application workflow, with red labels indicating approximate computation times: (1) An initial random structure (containing $\sim 500$ atoms) is generated in under 1 s; (2) this structure is passed to external tools (e.g., LAMMPS or first-principles calculations) for SRO evaluation (taking $\sim 25$ min in LAMMPS); (3) the resulting structure is analyzed in under 1 s to determine the target SRO; (4) a large-scale target structure (containing $256{,}000$ atoms) is constructed in about 35 s, based on the specified SRO parameters.
As illustrated in Fig. 3, PyHEA significantly streamlines the modeling workflow for HEAs. A small-scale system of approximately 500 atoms is first generated and refined (details of the Monte Carlo steps and convergence criteria are provided in section 4), after which the final SRO parameters are extracted. These SRO parameters guide the construction of a large-scale configuration containing up to $256{,}000$ atoms, a process completed in mere 35 seconds—substantially shorter than the hours or days typically required by Monte Carlo–Molecular Dynamics (MC–MD) simulations.
Figure 4: SRO values and Radial distribution function (RDF) for key atomic pairs at various temperatures in a $256{,}000$-atom Fe–Mn–Cr–Co system. (a) A radar plot illustrating SRO values for selected atomic pairs (e.g., Fe–Fe, Fe–Mn, Fe–Cr) at three temperatures: 750 K, 873 K, and 1373 K. (b) RDF curves showing the density of atomic pairs as a function of separation distance (Å) at the same temperatures.
To demonstrate this workflow, we generated a $256{,}000$-atom Fe–Mn–Cr–Co (40 % Fe, 40 % Mn, 10 % Cr, 10 % Co) system using PyHEA, then performed a 10 ps LAMMPS MD simulation in an NPT ensemble at multiple temperatures to obtain radial distribution function (RDF) data. Fig. 4 shows both the RDF and the SRO for representative atomic pairs. The RDF peaks indicate the positions of the first and second coordination shells, while the corresponding SRO values reveal how strongly specific atomic species tend to cluster or avoid one another.
The RDF and SRO curves exhibit a clear transition from a random-like distribution at 1373 K to increasingly pronounced SRO configurations at 750 K. This trend aligns qualitatively with prior computational and experimental studies on Fe–Mn–Cr–Co alloys, underscoring the physical reliability and robustness of our approach.
Finally, additional temperature-dependent SRO details are illustrated in Fig. 7, which visualizes the Warren–Cowley SRO parameters at 1373 K, 873 K, and 750 K. Taken together, these results underscore PyHEA’s efficiency in handling large-scale atomistic simulations and its capability to produce physically meaningful predictions that align well with established literature.